The vectors V1 - (1, 1, 2, 4), V2 - (2, - I , -5 , 2), V3 = (1, -1 , -4 , 0) and

V4 = (2, 1, 1 ,6 ) are linearly independent. Is it true ? Justify your answer


1
Expert's answer
2022-04-15T11:06:18-0400

Let "A =\begin{pmatrix}
 1 & 1 & 2 & 4\\
 2 & -1 & -5 & 2\\
 1 & -1 & -4 & 0\\
 2 & 1 & 1 & 6
\end{pmatrix}"

"\begin{pmatrix}
 1 & 1 & 2 & 4\\
 2 & -1 & -5 & 2\\
 1 & -1 & -4 & 0\\
 2 & 1 & 1 & 6
\end{pmatrix}" ~ "[V_2 -2V_1, V_3 - V_1, V_4 - 2V_1]" ~"\begin{pmatrix}
 1 & 1 & 2 & 4\\
 0 & -3 & -9 & -6\\
0 & -2 & -6 & -4 \\
0 & 0 & -1 & -2
\end{pmatrix}" ~

~"[V_3 - \cfrac{2}{3}V_2]" ~ "\begin{pmatrix}
 1 & 1 & 2 & 4\\
 0 & -3 & -9 & -6\\
0 & 0&0&0\\
0&0&-1&-2
\end{pmatrix}"

So we can see "rank \ A = 3 < 4"

Hence the vectors are linearly dependent